Circular No. 8749

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 S1 (CHRISTENSEN) E. J. Christensen reports the discovery of another comet on Catalina

Sky Survey images taken with the 0.68-m Schmidt telescope, the object displaying a well-condensed 15′′ coma and a 2′ straight and narrow tail in p.a. 245^{o} (discovery observation tabulated below). Following posting on the ‘NEO Confirmation Page’, numerous CCD observers have reported information on the comet’s coma diameter: Sept. 16.9 UT, 13′′ (J. Lacruz, Madrid, Spain, 0.30-m reflector); 17.3, 10′′ (J. Young, Table Mountain, 0.61-m reflector; with a bright, round central condensation); 17.3, 6′′ (J. E. McGaha, Tucson, AZ, 0.62-m reflector, six 60-s stacked images); 17.35– 17.38, ≥ 12′′ FWHM (D. T. Durig, T. E. Stegall, and E. A. Pierce, Sewanee, TN, 0.30-m reflector). Reported tail lengths and position angles: Sept. 16.9, 32′′, 235^{o} (Lacruz); 16.9, 3′, 237^{o} (A. Baransky, Kiev, 0.7-m reflector); 16.94, 25′′, 230^{o} (G. Sostero and E. Guido, Remanzacco, Italy, 0.25-m reflector); 16.95–16.96, –, 229^{o} (R. Ferrando and M. Ferrando, Valencia, Spain, 0.36-m reflector); 17.3, 80′′, 235^{o} (Young; tail curves slightly after 30′′ with a p.a. spanning 235^{o}–265^{o}, showing a slight fan shape); 17.3, 211′′ (10′′ wide), 217^{o} (McGaha); 17.35–17.38, 1′.5, ∼ 237^{o} (Durig et al.).

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.16.24646  23^{h}25^{m}47.^{s}64  +24^{o}50′12′′.9  17.4

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2006-S07.

T = 2006 Aug. 27.422 TT  ω = 124.^{o}815
e = 0.61999  Ω = 214.641 }2000.0
q = 1.35354 AU  i = 11.610
a = 3.56182 AU n^{o} = 0.146621 P = 6.7 years

SUPERNOVAE 2006ew–2006fn On CBETs 621 and 623, J. Frieman reports the discovery details for

18 supernovae on behalf of the Sloan Digital Sky Survey II collaboration. All the new objects are at g magnitude 20 or fainter except for 2006ez (predicted to reach mag 19.3 in mid-Sept.), 2006fd (reported at mag 19.4 on Aug. 27), and 2006fg (at mag 19.1 on Aug. 28). All are type-Ia events except for two type-II supernovae (2006ez and 2006fg) and 2006fe (whose type was not yet determined).

2006 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green